OpenStudy (anonymous):

Jonas cut a piece of rope into 6 pieces that were each 5 3/8 inches long. How long was the original piece of rope?

OpenStudy (anonymous):

no. of pieces = 6 length of a piece = 5(3/8) = 43/8 total length = 6* 43/ 8 = 139/4 = 34(3/4)

OpenStudy (anonymous):

what u mean?

OpenStudy (anonymous):

Jonas cut a piece of rope into 6 pieces And each piece was 5 3/8 inches long. so if 1 piece is 5(3/8) in. long and the rope was divided into 6 pieces then what wll be the length?

OpenStudy (ageta):

you just multiply 5 3/8 by 6